Check whether the given function is odd or even
f(x) = ln|(1-e⁻ˣ)/(1+e⁻ˣ)|

please its urgent answer it.
1
Expert's answer
2019-10-22T08:36:58-0400


We are going to check the given function is Odd (or) Even.



A function "f(X)" is said to be Odd function if

"{f(-X)} = - f(X)"

A function "f(X)" is said to be Even function if

"{f(-X)} = + f(X)"

The given function is

"f(X) = ln | \frac { 1 - e^{-X}} { 1 + e^{-X}}|"

Now plug X = - X in the function "f(X)":


"f(- X) = ln |\frac {1- e^ {-(-X)}} {1 + e^ {-(-X)}}|"


"f(- X) = ln |\frac {1- e^ {(+X)}} {1 + e^ {(+X)}}|"


"f(- X) = ln |\frac {1- e^ {(X)}} {1 + e^ {(X)}}|"

Now, divide the numerator and denominator by e^X



"f(- X) = ln |\frac {\frac {1} { e^ {X}} -1} {\frac {1} { e^ {X}} +1}|"

"f(- X) = ln |\frac { e^ {-X} -1} { { e^ {-X}} +1}|"

"f(- X) = ln |\frac { -(-e^{-X} +1)} { { e^ {-X}} +1}|"


"f(- X) = ln |\frac { (-e^ {-X} +1)} { { e^ {-X}} +1}|"

Since "| - a | = |a|",

"f(- X) = ln |\frac { (1 -e^ {-X})} { { 1 + e^ {-X}} }| = f(X)"

Thus, it is an Even function.


Answer: The given function is Even